OpenStudy (anonymous):

A geometric sequence is 3 4 , 9, 108, 1296, 15,552, 186,624, ... . Which is the general term of the sequence?

OpenStudy (anonymous):

a = 3 , d = 4/3 Tn = ar^ (n-1) = 3 x 4/3 ( n-1 ) ??? hope i help

OpenStudy (anonymous):

I'll post the answer choices.

OpenStudy (anonymous):

ohh its 3/4 ... i thought 3, 4, 108 etc...

OpenStudy (anonymous):

so you have a general formula for geometric sequences which is Tn = ar^(n-1) suppose you have got a sequence like : 2 , 4 , 6 , 8 , 10 we denote the first term with letter "a" which = 2 and r is the common ration and = 2nd term divide by 1st term = 4 divide by 2 = 2 therefore : Tn = 4^(n-1)

OpenStudy (anonymous):

ratio* in your case : a = 3/4 and r = 9 divide by 3/4 r=12 so general form : Tn = 3/4 ( 12) ^ (n-1)

OpenStudy (anonymous):

Thank you so much. :D